OpenStudy (anonymous):

(x-3)e^x=0 ? How to figure this out?

OpenStudy (turingtest):

by the zero factor property either e^x=0 or x-3=0 when is e^x=0 ?

OpenStudy (anonymous):

or @ -infi :P

OpenStudy (turingtest):

...and as we just established x=-infinity is not an actual possible value of x, so for all intents and purposes, when is e^x=0 (let the asker answer please)